OpenStudy (lgbasallote):

how to find the percent composition by mass of \(\text C_8 \text H_8\)

OpenStudy (lgbasallote):

would appreciate @waterineyes 's expertise

OpenStudy (anonymous):

I have to go now...

OpenStudy (lgbasallote):

that's convenient

OpenStudy (lgbasallote):

@mathslover are you a chemslover too?

mathslover (mathslover):

oh k .. so wait

mathslover (mathslover):

percent composition by mass of C and H both ?

mathslover (mathslover):

first of all: mass of C = 12 (approx.) H = 1 so mass of C_8 = 12*8 = 96 g and that of H_8 = 1* 8 = 8 g

mathslover (mathslover):

total mass = 104 g percent composition of mass of C in C_8 H_8 = \(\large{\frac{96}{104}*100 }\) = 92.3 % percent composition of mass of H in C_8 H_8 = \(\large{\frac{8}{104}*100}\) = 7.7 %

mathslover (mathslover):

got it?

OpenStudy (lgbasallote):

well i know the mass percent of the elements...what intrigues me is the mass % of the compound

mathslover (mathslover):

% composition by mass of a compound means : mass % of the element in the compound..

OpenStudy (lgbasallote):

just that?

mathslover (mathslover):

though .. @Xishem can confirm me .. if you want confirmation ..

mathslover (mathslover):

may be m wrong somewhere

OpenStudy (lgbasallote):

interestiing

OpenStudy (lgbasallote):

it seems there's more to this

OpenStudy (xishem):

What mathslover said seems accurate. It's as simple as converting from moles of each element to mass of each element and dividing each part of the mass by the total mass. The percent composition of the compound IS the percent of carbon by mass as well as the percent of hydrogen by mass. I'm unsure what you are uncertain about.

OpenStudy (lgbasallote):

well what i am unsure about is this http://openstudy.com/updates/5039df93e4b043c156a326fd idk if you really use this method in that question...or it was just a coincidence